Feed, Tiger, and Nancy play a round of golf together. Their combined score is 223. Fred's score was 3 more than Nancy's, and Tiger's score was 7 more than Fred's. What was each person's score?
Answer by ewatrrr(24785)


Hi,
Let x repesent Nancy's score, then Fred's is (x+3) and Tiger's is (x+3+7) or (x+10)
x + (x + 3) + (x + 10) = 223
solve for x
3x + 13 = 223
3x = 210
x = 70
Nancy's score was 70, Fred's 73 and Tiger's 80.
checking our answer
70 + 73 + 80 = 223